OpenStudy (susanna):

Can someone check my answer? Simplify the expression cos x + sin x tan x My answer: sec x

OpenStudy (anonymous):

That's correct.

OpenStudy (susanna):

Thanx!

OpenStudy (susanna):

It was kind of a guess, so can you please explain how to get that answer?

OpenStudy (susanna):

@mathmale @mathstudent55 Please HELP

OpenStudy (susanna):

@SolomonZelman

OpenStudy (anonymous):

Sure. Here's how I would do it (there are multiple methods I'm sure): First: Convert tan(x) to sin and cos\[\cos(x)+\sin(x)(\frac{ \sin(x) }{ \cos(x) })\] Now multiply the sin(x) \[\cos(x)+\frac{ \sin ^{2}(x) }{ \cos(x) }\] Then find a common denominator \[\cos(x)(\frac{ \cos(x) }{ \cos(x) })+\frac{ \sin ^{2}(x) }{ \cos(x) }\] Multiply the cos(x) and make it all one fraction\[\frac{ \cos ^{2}(x)+\sin ^{2}(x) }{ \cos(x) }\] Now by Trig identities you know that\[\sin ^{2}(x)+\cos ^{2}(x)=1\] So it becomes \[\frac{ 1 }{ \cos(x) }=\sec(x)\]

OpenStudy (susanna):

Wow, thank you sooo much. Now I get it.

OpenStudy (anonymous):

You're welcome, I'm glad I could help :)